OpenStudy (anonymous):

Simplify. -14x^3/x^3-5x^4

OpenStudy (whpalmer4):

Do you see any common factors you can remove from numerator and denominator?

sam (.sam.):

well if you factor x^3 from the denominator then you'll cancel out the x^3 with the numerator

OpenStudy (whpalmer4):

-1 can also be factored out of both

OpenStudy (anonymous):

ok so -13x^3/x^3-4x^4 ??

OpenStudy (whpalmer4):

Not subtracted, factored. \[\frac{-14x^3}{x^3-5x^4} = \frac{\cancel{x^3}(-14)}{\cancel{x^3}(1-5x)}\]

OpenStudy (whpalmer4):

\[\frac{-14x^3}{x^3-5x^4} = \frac{\cancel{x^3}(-14)}{\cancel{x^3}(1-5x )} = \frac{\cancel{-1}(14)}{\cancel{-1}(5x-1)}=\frac{14}{5x-1}\]